题目
A、B 为两事件,若 P(A cup B) = 0.8,P(A) = 0.2,P(overline(B)) = 0.4,则(A) P(overline(A) overline(B)) = 0.32(B) P(overline(A) overline(B)) = 0.2(C) P(B - A) = 0.4(D) P(overline(B) A) = 0.48
$A$、$B$ 为两事件,若 $P(A \cup B) = 0.8$,$P(A) = 0.2$,$P(\overline{B}) = 0.4$,则
(A) $P(\overline{A} \overline{B}) = 0.32$
(B) $P(\overline{A} \overline{B}) = 0.2$
(C) $P(B - A) = 0.4$
(D) $P(\overline{B} A) = 0.48$
题目解答
答案
已知条件:
- $P(A \cup B) = 0.8$
- $P(A) = 0.2$
- $P(\bar{B}) = 0.4$,则 $P(B) = 1 - P(\bar{B}) = 0.6$
计算 $P(A \cap B)$:
$P(A \cup B) = P(A) + P(B) - P(A \cap B) \implies 0.8 = 0.2 + 0.6 - P(A \cap B) \implies P(A \cap B) = 0$
事件 $A$ 和 $B$ 互斥。
选项分析:
(A) $P(\bar{A} \cap \bar{B}) = P(\overline{A \cup B}) = 1 - P(A \cup B) = 1 - 0.8 = 0.2$,错误。
(B) $P(\bar{A} \cap \bar{B}) = 0.2$,正确。
(C) $P(B - A) = P(B \cap \bar{A}) = P(B) - P(A \cap B) = 0.6 - 0 = 0.6$,错误。
(D) $P(\bar{B} \mid A) = \frac{P(\bar{B} \cap A)}{P(A)} = \frac{P(A) - P(A \cap B)}{P(A)} = \frac{0.2}{0.2} = 1$,错误。
答案: $\boxed{B}$